Yellow Sand Banking is evaluating a project that would require the purchase of a piece of equipment for 790,000 dollars today. During year 1, the project is expected to have relevant revenue of 747,000 dollars, relevant costs of 300,200 dollars, and relevant depreciation of 50,100 dollars. Yellow Sand Banking would need to borrow 790,000 today for the equipment and would need to make an interest payment of 33,800 dollars to the bank in 1 year. Relevant operating cash flow for the project in year 1 is expected to be 226,473 dollars. What is the tax rate expected to be in year 1? Answer as a rate in decimal format so that 12.34% would be entered as .1234 and 0.98% would be entered as .0098.
Answer #1
OCF 226473
Less: Depreciation 50100
Net Income 176373
Revenue 747000
Less: Costs 300200
Depreciation 50100
Interest 33800
Profit before tax 362,900

Tax = PBT-Net Income = 362900-176373 = 186527

Tax rate = 186527/362900 = 0.5140